Two scales on a nondigital voltmeter measure voltages up to 20.0 and 27.0 V, respectively. The resistance connected in series with the galvanometer is 1680 Ω for the 20.0-V scale and 2930 Ω for the 27.0-V scale. Determine the coil resistance.
Answer #1

let the coil resistance is r

let the full scale current is I

for measuring 20 V

at full scale

20 = I *(1680 + r) ---(1)

for 27 V

27 = I*(2930 + r) ---(2)

from equation 1 and 2

27/20 =(2930 + r)/ (1680 + r)

solving for r

r = 1891.4 Ohm

the coil resistance is 1891.4 Ohm